# Enhanced Rényi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees

### v1

@@ Abstract
This paper presents an enhanced post-quantum key agreement protocol based on Rényi entropy, addressing vulnerabilities in the original construction while preserving information-theoretic security properties. We develop a theoretical framework leveraging entropy-preserving operations and secret-shared verification to achieve provable security against quantum adversaries. Through entropy amplification techniques and quantum-resistant commitments, the protocol establishes $2^{128}$ quantum security guarantees under the quantum random oracle model. Key innovations include a confidentiality-preserving verification mechanism using distributed polynomial commitments, tightened min-entropy bounds with guaranteed non-negativity, and composable security proofs in the quantum universal composability framework. Unlike computational approaches, our method provides information-theoretic security without hardness assumptions while maintaining polynomial complexity. Theoretical analysis demonstrates resilience against known quantum attack vectors, including Grover-accelerated brute force and quantum memory attacks. The protocol achieves parameterization for 128-bit quantum security with efficient $\mathcal{O}(n^{2})$ communication complexity. Extensions to secure multiparty computation and quantum network applications are established, providing a foundation for long-term cryptographic security. All security claims are derived from mathematical proofs; this theoretical work presents no experimental validation.
Index Terms:
Rényi entropy, post-quantum cryptography, key agreement, information-theoretic security, quantum security, entropy preservation, secret sharing, quantum random oracle model.
@@ 1 Introduction
The anticipated development of large-scale quantum computing represents a significant challenge to contemporary cryptographic infrastructure [31, 20, 14, 1]. Shor’s polynomial-time quantum algorithm for integer factorization and discrete logarithms [1] compromises the security of widely-deployed asymmetric cryptosystems including RSA, ECC, and Diffie-Hellman key exchange. Grover’s quadratic speedup for unstructured search [2] reduces the effective security strength of symmetric primitives, necessitating key size increases [27]. This quantum threat landscape has catalyzed post-quantum cryptography (PQC) development, with NIST standardizing lattice-based, code-based, and multivariate schemes [20, 9]. However, these approaches rely on computational hardness assumptions that may be vulnerable to algorithmic advances [26, 28, 10].
1-A Research Context and Motivation
@@
The original Rényi entropy-based key agreement protocol [32] demonstrated theoretical promise but contained vulnerabilities: (1) input exposure during broadcast enabling key compromise, (2) potentially negative entropy bounds violating security requirements, and (3) insufficient protection against quantum-specific attacks. This work addresses these limitations through three key innovations:
@@
Theorem 1 (Quantum Entropy Preservation Bound).
@@ 1-B Key Contributions
1. Confidentiality-Preserving Verification: A secret-shared verification mechanism using distributed polynomial commitments that prevents input exposure while enabling secure entropy verification, addressing a vulnerability in prior constructions [32].
2. Optimal Entropy Bounds: Mathematically rigorous min-entropy preservation theorems with guaranteed non-negative bounds and parameterization strategies for quantum security.
3. Quantum Universal Composability: Composable security proof for information-theoretic key agreement in the quantum universal composability (QUC) model [5, 6], demonstrating secure realization of ideal key functionality.
4. Quantum Attack Resistance: Formal analysis against known quantum attack vectors including Grover-accelerated brute force, quantum collision search [7], and quantum memory attacks [21], with quantifiable security bounds.
5. Entropy Amplification Framework: Generalized framework for multi-party entropy amplification using R’enyi entropy measures [3], providing exponential security scaling against quantum adversaries.
6. Hybrid Security Extension: Integration with quantum key distribution (QKD) [25] that enhances security against active adversaries while preserving information-theoretic guarantees.
7. Secure Computation Extension: Secure extension to general secure multiparty computation [24, 17] for linear functions, enabling privacy-preserving applications.
@@
1-C Quantum Entropic Framework
The foundation of our approach rests upon the rigorous quantification of uncertainty in quantum systems. Unlike classical entropy measures, quantum Rényi entropy (9) captures the fundamental limits of information extraction under quantum mechanical constraints. This becomes crucial when adversaries possess quantum memory capable of storing superposition states for delayed measurement [21]. Our framework explicitly addresses this quantum advantage by establishing composable entropy bounds through:
@@ 2 Background and Theoretical Foundations
2-A Quantum Computing Threat Model
@@
Definition 1 (Quantum Polynomial-Time Adversary).
A quantum adversary $\mathscr{A}$ is a polynomial-time quantum algorithm with:
@@
  • Brassard-Høyer-Tapp (BHT): Quantum collision finding in $\mathcal{O}(2^{m/3})$ time [7]
@@
2-B Rényi Entropy Framework
Rényi entropy [3] provides a parametric family of entropy measures essential for cryptographic security analysis. Its operational significance in quantum cryptography stems from connections to guessing probabilities [18]:
Definition 2 (Rényi Entropy).
For discrete random variable $X\sim P_{X}$ over $\mathcal{X}$, Rényi entropy of order $\alpha$ is:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\left(\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}\right)$ (2)
@@ 2-B Rényi Entropy Framework
$\lim_{\alpha\to 1}H_{\alpha}(X)=H(X)=-\sum_{x}P_{X}(x)\log P_{X}(x)$ (Shannon) (3)
$H_{2}(X)=-\log\left(\sum_{x}P_{X}(x)^{2}\right)$ (Collision) (4)
$\lim_{\alpha\to\infty}H_{\alpha}(X)=H_{\infty}(X)=-\log\max_{x}P_{X}(x)$ (Min-entropy) (5)
@@
Lemma 1 (Entropy Transformation).
@@
2-C Quantum Information Theoretic Foundations
Definition 3 (Classical vs. Quantum Entropy Notation).
@@ 2-C Quantum Information Theoretic Foundations
  • $H_{\alpha}(X)$ denotes classical Rényi entropy for random variable $X$
  • $S_{\alpha}(\rho)$ denotes quantum Rényi entropy for density operator $\rho$
@@
Definition 4 (Quantum Rényi Entropy).
For density operator $\rho$, quantum Rényi entropy is:
$S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{tr}(\rho^{\alpha})$ (9)
@@
  • Quantum collision entropy: $S_{2}(\rho)=-\log\operatorname{tr}(\rho^{2})$ quantifies state distinguishability
@@
2-D Distributed Verification via Secret Sharing
@@
Theorem 2 (Shamir’s Secret Sharing).
@@
2-E Symbol Table
TABLE 1: Mathematical Symbols and Definitions
Symbol Meaning
$n$ Number of protocol participants
$\kappa$ Security parameter (bit strength)
$m$ Bit-length of secret shares
$\gamma$ Min-entropy threshold per share
$t$ Reconstruction threshold ($t=\lfloor n/2\rfloor+1$)
$\delta$ Entropy estimation accuracy
$H_{\alpha}(X)$ Classical Rényi entropy of order $\alpha$
$S_{\alpha}(\rho)$ Quantum Rényi entropy
$H_{\infty}(X)$ Min-entropy: $\log\max_{x}P_{X}(x)$
$\hat{H}_{i}$ Estimated min-entropy of $s_{i}$
$s_{i}$ Secret share generated by party $i$
$c_{i}$ Cryptographic commitment: $\mathcal{H}(s_{i}\|\hat{H}_{i})$
$f_{i}(x)$ Sharing polynomial for $s_{i}$
$sh_{ij}$ Share of $s_{i}$ distributed to $P_{j}$
$\mathbb{F}_{2^{m}}$ Finite field with $2^{m}$ elements
$\mathcal{H}$ Quantum random oracle
$\oplus$ Bitwise XOR operation
$Q$ Adversary quantum memory bound ($Q=2^{\kappa/2}$ qubits)
$\mathscr{A}$ Quantum polynomial-time adversary
$\mathcal{F}_{KA}$ Ideal key agreement functionality
$K$ Established secret key
$\lambda$ Commitment security parameter ($\lambda=2\kappa$)
$\epsilon$ Security error bound ($\epsilon=2^{-40}$)
@@ 3 Enhanced Protocol Design
3-A System Model and Threat Analysis
Consider a network of $n$ parties $\mathcal{P}=\{P_{1},\dots,P_{n}\}$ establishing shared secret key $K\in\{0,1\}^{\kappa}$.
@@ 3-A System Model and Threat Analysis
  • Quantum random oracle access to $\mathcal{H}$
@@
3-B Confidentiality-Preserving Verification Mechanism
@@
Definition 5 (Distributed Polynomial Commitment).
@@
with $a_{k}\xleftarrow{\$}\{0,1\}^{m}$. The commitment is $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ where $\hat{H}_{i}\approx H_{\infty}(s_{i})$.
@@
    Verifies $c_{i}=\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})$
@@
Theorem 3 (Verification Security).
@@
$\Pr[\text{successful verification}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}$ (12)
@@ 3-C Key Agreement Protocol
Algorithm 1 Enhanced Rényi Key Agreement Protocol
0: Security parameter $\kappa$, min-entropy threshold $\gamma$, parties $n$, estimation accuracy $\delta$
0: Shared secret key $K$
 Set $t=\lceil n/2\rceil+1$ {Honest majority threshold}
Phase 1: Initialization
for each party $P_{i}\in\mathcal{P}$ do
  Sample $s_{i}\leftarrow\{0,1\}^{m}$ with $H_{\infty}(s_{i})\geq\gamma$ {High-entropy secret}
  Estimate $\hat{H}_{i}$ such that $|H_{\infty}(s_{i})-\hat{H}_{i}|\leq\delta$ {Min-entropy estimation}
  Compute commitment $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ {QROM-based binding}
  Broadcast $c_{i}$ to all parties {Public commitment}
end for
Phase 2: Share Distribution
for each $P_{i}\in\mathcal{P}$ do
  Generate random polynomial $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}$, $a_{k}\xleftarrow{\$}\{0,1\}^{m}$
  for each $P_{j}\in\mathcal{P}\setminus\{P_{i}\}$ do
   Compute share $sh_{ij}=f_{i}(j)$
   Securely send $(i,sh_{ij})$ to $P_{j}$ via authenticated channel
  end for
end for
Phase 3: Verification
for each $P_{j}\in\mathcal{P}$ do
  for each $P_{i}\in\mathcal{P}\setminus\{P_{j}\}$ do
   Collect $\geq t$ valid shares $\{sh_{ik}\}_{k\in S}$ for $|S|\geq t$
   Reconstruct $\tilde{s}_{i}=\sum_{k\in S}sh_{ik}\cdot L_{k}(0)$ {Lagrange interpolation}
   Verify $\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})=c_{i}$ {Commitment consistency}
   Verify $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ {Entropy threshold}
   if any verification fails then
    Abort protocol and output $\bot$
   end if
  end for
end for
Phase 4: Key Derivation
 Each $P_{i}$ reveals $s_{i}$ to all parties via authenticated channels {Safe after verification}
 Compute $S=\bigoplus_{i=1}^{n}s_{i}$ {Entropy-preserving combination}
 Compute $K=\mathcal{H}(S)$ {Quantum-secure randomness extraction}
return  $K$
@@
3-D Entropy Verification Theoretical Basis
@@
    $\hat{H}_{\infty}=-\log\left(\max_{x\in\mathcal{S}}\hat{P}(x)\right)\pm\delta$ (13)
@@
3-E Quantum Security Design Principles
@@
    Superposition Resistance: Polynomial commitments use injective encoding $\mathcal{H}(s_{i}\parallel\hat{H}_{i})$ to prevent quantum ambiguity attacks. The concatenation ensures $\forall s_{i}\neq s_{j},\mathcal{H}(s_{i}\parallel\cdot)\cap\mathcal{H}(s_{j}\parallel\cdot)=\emptyset$ with probability $1-\operatorname{negl}(\lambda)$ under QROM.
@@
[…]
    $\Delta(\rho_{SE},\rho_{S}\otimes\rho_{E})\leq 2^{-H_{\min}(S|E)}$ (14)
@@
    Commit $\rightarrow\text{Share}\rightarrow\text{Verify}$
    $\Downarrow$
    No-cloning theorem enforced

### v2

# Enhanced Renyi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees

@@ Abstract
This paper presents an enhanced post-quantum key agreement protocol based on Renyi entropy, addressing vulnerabilities in the original construction while preserving information-theoretic security properties. We develop a theoretical framework leveraging entropy-preserving operations and secret-shared verification to achieve provable security against quantum adversaries. Through entropy amplification techniques and quantum-resistant commitments, the protocol establishes $2^{128}$ quantum security guarantees under the quantum random oracle model. Key innovations include a confidentiality-preserving verification mechanism using distributed polynomial commitments, tightened min-entropy bounds with guaranteed non-negativity, and composable security proofs in the quantum universal composability framework. Unlike computational approaches, our method provides information-theoretic security without hardness assumptions while maintaining polynomial complexity. Theoretical analysis demonstrates resilience against known quantum attack vectors, including Grover-accelerated brute force and quantum memory attacks. The protocol achieves parameterization for 128-bit quantum security with efficient $\mathcal{O}(n^{2})$ communication complexity. Extensions to secure multiparty computation and quantum network applications are established, providing a foundation for long-term cryptographic security.
keywords:
Renyi entropy , post-quantum cryptography , key agreement , information-theoretic security , quantum security , entropy preservation , secret sharing , quantum random oracle model
journal: Theoretical Computer Science
[…]
organization=Independent Researcher,city=,country=
[…]
organization=School of Electrical and Computer Engineering, Nanfang College Guangzhou,addressline=, city=Guangzhou, postcode=510970, country=China
@@ 1 Introduction
The anticipated development of large-scale quantum computing represents a significant challenge to contemporary cryptographic infrastructure [1, 20, 14]. Shor’s polynomial-time quantum algorithm for integer factorization and discrete logarithms [1] compromises the security of widely-deployed asymmetric cryptosystems including RSA, ECC, and Diffie-Hellman key exchange. Grover’s quadratic speedup for unstructured search [2] reduces the effective security strength of symmetric primitives, necessitating key size increases [27]. This quantum threat landscape has catalyzed post-quantum cryptography (PQC) development, with NIST standardizing lattice-based, code-based, and multivariate schemes [20, 9]. However, these approaches rely on computational hardness assumptions that may be vulnerable to algorithmic advances [10, 28].
1.1 Research Context and Motivation
@@
The original Renyi entropy-based key agreement protocol [32] demonstrated theoretical promise but contained vulnerabilities: (1) input exposure during broadcast enabling key compromise, (2) potentially negative entropy bounds violating security requirements, and (3) insufficient protection against quantum-specific attacks. This work addresses these limitations through three key innovations:
@@
Theorem 1.1 (Quantum Entropy Preservation Bound).
@@ 1.2 Key Contributions
  1. 1.
    Confidentiality-Preserving Verification: A secret-shared verification mechanism using distributed polynomial commitments that prevents input exposure while enabling secure entropy verification, addressing a vulnerability in prior constructions [32].
  2. 2.
    Optimal Entropy Bounds: Mathematically rigorous min-entropy preservation theorems with guaranteed non-negative bounds and parameterization strategies for quantum security.
  3. 3.
    Quantum Universal Composability: Composable security proof for information-theoretic key agreement in the quantum universal composability (QUC) model [5, 6], demonstrating secure realization of ideal key functionality.
  4. 4.
    Quantum Attack Resistance: Formal analysis against known quantum attack vectors including Grover-accelerated brute force, quantum collision search [7], and quantum memory attacks [21], with quantifiable security bounds.
  5. 5.
    Entropy Amplification Framework: Generalized framework for multi-party entropy amplification using R’enyi entropy measures [3], providing exponential security scaling against quantum adversaries.
  6. 6.
    Hybrid Security Extension: Integration with quantum key distribution (QKD) [25] that enhances security against active adversaries while preserving information-theoretic guarantees.
  7. 7.
    Secure Computation Extension: Secure extension to general secure multiparty computation [17, 24] for linear functions, enabling privacy-preserving applications.
@@
1.3 Quantum Entropic Framework
The foundation of our approach rests upon the rigorous quantification of uncertainty in quantum systems. Unlike classical entropy measures, quantum Renyi entropy (4) captures the fundamental limits of information extraction under quantum mechanical constraints. This becomes crucial when adversaries possess quantum memory capable of storing superposition states for delayed measurement [21]. Our framework explicitly addresses this quantum advantage by establishing composable entropy bounds through:
  • 1.
@@ 2 Background and Theoretical Foundations
2.1 Quantum Computing Threat Model
@@
Definition 2.1 (Quantum Polynomial-Time Adversary).
A quantum adversary $\mathcal{A}$ is a polynomial-time quantum algorithm with:
  • 1.
@@
  • 3.
    Brassard-Hoyer-Tapp (BHT): Quantum collision finding in $\mathcal{O}(2^{m/3})$ time [7]
  • 4.
@@
2.2 Renyi Entropy Framework
Renyi entropy [3] provides a parametric family of entropy measures essential for cryptographic security analysis. Its operational significance in quantum cryptography stems from connections to guessing probabilities [18]:
Definition 2.2 (Renyi Entropy).
For discrete random variable $X\sim P_{X}$ over $\mathcal{X}$, Renyi entropy of order $\alpha$ is:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\left(\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}\right)$
@@ 2.2 Renyi Entropy Framework
$\lim_{\alpha\to 1}H_{\alpha}(X)$ $=H(X)=-\sum_{x}P_{X}(x)\log P_{X}(x)\quad\text{(Shannon)}$
$H_{2}(X)$ $=-\log\left(\sum_{x}P_{X}(x)^{2}\right)\quad\text{(Collision)}$
$\lim_{\alpha\rightarrow\infty}H_{\alpha}(X)$ $=H_{\infty}(X)=-\log\max_{x}P_{X}(x)\quad\text{(Min-entropy)}$
@@
Lemma 2.3 (Entropy Transformation).
@@
2.3 Quantum Information Theoretic Foundations
Definition 2.4 (Classical vs. Quantum Entropy Notation).
@@ 2.3 Quantum Information Theoretic Foundations
  • 1.
    $H_{\alpha}(X)$ denotes classical Renyi entropy for random variable $X$
  • 2.
    $S_{\alpha}(\rho)$ denotes quantum Renyi entropy for density operator $\rho$
  • 3.
@@
Definition 2.5 (Quantum Renyi Entropy).
For density operator $\rho$, quantum Renyi entropy is:
$S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\textnormal{tr}(\rho^{\alpha})$ (4)
@@
  • 2.
    Quantum collision entropy: $S_{2}(\rho)=-\log\textnormal{tr}(\rho^{2})$ quantifies state distinguishability
  • 3.
@@
2.4 Distributed Verification via Secret Sharing
@@
Theorem 2.6 (Shamir’s Secret Sharing).
@@ 3 Enhanced Protocol Design
3.1 System Model and Threat Analysis
Consider a network of $n$ parties $\mathcal{P}=\{P_{1},\ldots,P_{n}\}$ establishing shared secret key $K\in\{0,1\}^{\kappa}$.
@@ 3.1 System Model and Threat Analysis
  • 2.
    Quantum random oracle access to hash functions
  • 3.
@@
3.2 Confidentiality-Preserving Verification Mechanism
@@
Definition 3.1 (Distributed Polynomial Commitment).
@@
with $a_{k}\overset{\mathdollar}{\leftarrow}\{0,1\}^{m}$. The commitment is $c_{i}=\mathcal{H}(s_{i}||\hat{H}_{i})$ where $\hat{H}_{i}\approx H_{\infty}(s_{i})$.
@@
    Verifies $c_{i}=\mathcal{H}(\tilde{s}_{i}||\hat{H}_{i})$
@@
Theorem 3.2 (Verification Security).
@@
$\Pr[\text{successful verification}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}$
@@ 3.3 Key Agreement Protocol
Algorithm 1 Enhanced Renyi Key Agreement Protocol
1:Security parameter $\kappa$, min-entropy threshold $\gamma$, parties $n$, estimation accuracy $\delta$
2:Shared secret key $K$
3:Set $t=\lceil n/2\rceil+1$ $\triangleright$ Honest majority threshold
4:
5:Phase 1: Initialization
6:for each party $P_{i}\in\mathcal{P}$ do
7:  Sample $s_{i}\leftarrow\{0,1\}^{m}$ with $H_{\infty}(s_{i})\geq\gamma$ $\triangleright$ High-entropy secret
8:  Estimate $\hat{H}_{i}$ such that $|H_{\infty}(s_{i})-\hat{H}_{i}|\leq\delta$ $\triangleright$ Min-entropy estimation
9:  Compute commitment $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ $\triangleright$ QROM-based binding
10:  Broadcast $c_{i}$ to all parties $\triangleright$ Public commitment
11:end for
12:
13:Phase 2: Share Distribution
14:for each party $P_{i}\in\mathcal{P}$ do
15:  Generate random polynomial $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}$, $a_{k}\overset{\mathdollar}{\leftarrow}\{0,1\}^{m}$
16:  for each $P_{j}\in\mathcal{P}\setminus\{P_{i}\}$ do
17:    Compute share $sh_{ij}=f_{i}(j)$
18:    Securely send $(i,sh_{ij})$ to $P_{j}$ via authenticated channel
19:  end for
20:end for
21:
22:Phase 3: Verification
23:for each party $P_{j}\in\mathcal{P}$ do
24:  for each $P_{i}\in\mathcal{P}\setminus\{P_{j}\}$ do
25:    Collect $\geq t$ valid shares $\{sh_{ik}\}_{k\in S}$ for $|S|\geq t$
26:    Reconstruct $\tilde{s}_{i}=\sum_{k\in S}sh_{ik}\cdot L_{k}(0)$ $\triangleright$ Lagrange interpolation
27:    Verify $\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})=c_{i}$ $\triangleright$ Commitment consistency
28:    Verify $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ $\triangleright$ Entropy threshold
29:    if any verification fails then
30:     Abort protocol and output $\bot$
31:    end if
32:  end for
33:end for
34:
35:Phase 4: Key Derivation
36:Each $P_{i}$ reveals $s_{i}$ to all parties via authenticated channels $\triangleright$ Safe after verification
37:Compute $S=\bigoplus_{i=1}^{n}s_{i}$ $\triangleright$ Entropy-preserving combination
38:Compute $K=\mathcal{H}(S)$ $\triangleright$ Quantum-secure randomness extraction
39:return $K$
@@
3.4 Entropy Verification Theoretical Basis
@@
    $\hat{H}_{\infty}=-\log\left(\max_{x\in\mathcal{S}}\hat{P}(x)\right)\pm\delta$
@@
3.5 Quantum Security Design Principles
@@
    Superposition Resistance: Polynomial commitments use injective encoding $\mathcal{H}(s_{i}\|\hat{H}_{i})$ to prevent quantum ambiguity attacks. The concatenation ensures $\forall s_{i}\neq s_{j},\mathcal{H}(s_{i}\|\cdot)\cap\mathcal{H}(s_{j}\|\cdot)=\emptyset$ with probability $1-\mathrm{negl}(\lambda)$ under QROM.
@@
    Entanglement Breaking: The final XOR operation $S=\bigoplus s_{i}$ acts as a non-commutative operator relative to quantum adversaries’ observation basis. For any quantum state $|\psi\rangle=\sum\alpha_{x,y}|x,y\rangle_{E}$, the operation satisfies:
    $\Delta(\rho_{SE},\rho_{S}\otimes\rho_{E})\leq 2^{-H_{\min}(S|E)}$
@@
    $\text{Commit }\rightarrow\text{Share }\rightarrow\text{Verify}$